OpenStudy (anonymous):

Quick question. Does 9sqrt(x) * 4x+1 not equal 36x+9sqrt(x)?

OpenStudy (anonymous):

Hello all.

OpenStudy (anonymous):

Any ideas?

OpenStudy (lukebluefive):

If you mean the following equation: \[9 \times \sqrt{x} \times 4x + 1 = 36x + 9 \times \sqrt{x}\] Then for very specific values of x, yes, but not in general.

OpenStudy (anonymous):

Then...what did I do wrong?

OpenStudy (lukebluefive):

Taking the left side of the equation: \[9 \times \sqrt{x} \times 4x + 1 = 36x \times \sqrt{x} + 1\] This is clearly not equal to the right. What did you start with?

OpenStudy (anonymous):

I mean...I started with the expression I mentioned in my question.

OpenStudy (lukebluefive):

I think you changed a multiply into an add and forgot you'd already multiplied by 9, so: \[36x \times \sqrt{x}\] Became: \[36x + 9 \times \sqrt{x}\]